A compound is found to contain 15.12 % phosphorus6.841 % nitrogen, and 78.03 % bromine by mass.


1
Expert's answer
2021-08-09T07:47:43-0400

Solution


P= 15.21%

N=6.841%

B=78.03%

Molar mass P=0.4881,N=0.4884,B=0.97

divide by the smallest number

"P=\frac{0.4881}{0.4881}=1"


"N=\frac{0.4884}{0.4881}=1"


"B= \frac{0.9765}{0.4881}= 2"


Empirical formula Br2NP